Nandini’s house is 9 / 10 km from her school. She walked some distance and then took a bus for 1 / 2 km to reach the school. How far did she walk?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Distance of the school from house = 9 / 10 km
Distance she travelled by bus = 1 / 2 km
Distance walked by Nandini = Total distance of the school–Distance she travelled by bus
= 9 / 10–1 / 2
= [(9× 1)–(1 × 5)] / 10
= (9–5) / 10
= 4 / 10
= 2 / 5 km
∴Distance walked by Nandini is 2 / 5 km

Thanks